5. The standard SI units of mass, time and length respactivey are a)grams, seconds, meters c) meters, seconds, kilograms b) kilograns, seconds, meters d) shugs, seconds, feet 6. A square piece of aluminium with as mass of 5.4 g has a length of 2 cm and a thickness of 5 mm. What is the density of the piece of aluminiun in 7. A ball is thrown into the air with a velocity of Sm/s from on top of a 36 m building. How long does it take the ball to hit the ground? 8. The velocity of a particle is v() (2-21) m/si+ (6t+2) m/sj What is the speed of the particle at t 3 s?
Answer #1

5.

SI unit of mass = kilograms = kg

SI unit of time = second

SI unit of length = meter

Correct option is B.

6.

Density is given by:

density = mass/Volume = m/V

Volume of aluminum wire = length*width*thickness = L*w*H

m = 5.4 gm = 5.4*10^-3 kg

L = w = 2 cm = 0.02 m

H = 5 mm = 0.005 m

So,

density = (5.4*10^-3)/(0.02*0.02*0.005) = 2700 kg/m^3

7.

Using 2nd kinematic equation

H = U*t + 0.5*a*t^2

H = -36 m

U = 8 m/sec

a = acceleration = -g = -9.8 m/sec^2

So,

-36 = 8*t + 0.5*(-9.8)*t^2

4.9*t^2 - 8*t - 36 = 0

Solving above quadratic equation

t = [8 + sqrt (8^2 + 4*36*4.9)]/(2*4.9)

t = 3.65 sec = time taken by ball to hit the ground

8.

V(t) = (t^2 - 2t) i + (6t + 2) j

At t = 3 sec

V(3) = (3^2 - 2*3) i + (6*3 + 2) j

V(3) = 3 i + 20 j

Speed = |V(3)|

Speed = sqrt (3^2 + 20^2) = 20.22 m/sec

Please Upvote.